OpenStudy (anonymous):

Write the inequality in SLOPE INTERCEPT FORM #2 9x+3y≥-12

OpenStudy (anonymous):

y>=3x-4

OpenStudy (johnweldon1993):

same thing here...we want to isolate y so 9x + 3y >= -12 subtract 9x from both sides 3y >= -12 - 9x divide both sides by 3 y >= 4 - 3x written correctly \[y \ge-3x + 4\]

OpenStudy (johnweldon1993):

sorry mistake.... 3y >= -12 - 9x divide both sides by 3 y >= -4 - 3x written correctly \[y \ge-3x - 4\]